# intrinsic motivation; active inference; empowerment; entropy

### v1

@@ 3.2 Constrained maximum occupancy
Thus, identifying a lack of action availability with a low-entropy state is plausible only in toy scenarios in which the entropy increase induced within the overall system by the dissolution of the agent is ignored. Death ought to be attractive to MOP agents unless they possess an a priori distinction between agent and environment, i.e. a “sense of self”. The upshot is that the implicit constraint enabling the emergence of goal-directed behavior in such agents is, in the general case, not simply long-term entropy maximization but also the existence of an agent with a repertoire of actions, encoded in the very partitioning of the space into action and state variables. Effectively, this amounts to a version of the “controllability” constraints that appear explicitly in active inference and empowerment, as the agent must exert control sufficient to enable homeostasis (i.e. the maintenance of internal states against dissipative forces). [^13] We may note that the ability to predict the entropy of future states so as to compute state value—however this is implemented—also corresponds to some local disequilibrium and thus imposes a de facto constraint on entropy (here, we are again explicitly considering the entropy of the internal states of the agent).
@@ 3.3 Model evidence and the will to live
Despite the arguments just given, the inversion of traditional assumptions about the relationship between exploration and reward highlighted by the MOP is appealing, as entropy maximization appears to be an essential feature of intelligence and life (England, 2013; Wissner-Gross and Freer, 2013; Sakthivadivel, 2022), more constant across distinct forms of life than any particular reward-seeking behavior. The idea that future path occupancy, as measured by entropy (Ramirez-Ruiz et al., 2022), is tantamount to remaining alive is one way of understanding the place of entropy maximization at the heart of accounts of intrinsic motivation.
We have seen however that in order to reproduce the goal-oriented behavior characteristic of complex biological intelligence, it is necessary to maximize entropy under the constraint that the agent’s existence, operationalized as a conditional independence between internal and external states (Friston, 2019), is maintained. Taking a page from Schopenhauer (Schopenhauer and Payne, 1958), intrinsic motivation may then be cast as simply the “will to live”, i.e. to persist as a living (moving, changing) thing, a basal impulse that takes different particular forms depending on local constraints (generative models). These constraints shape the primary motivational force of entropy production, such that conditional independence structures are maintained, effectively a (sometimes complex) form of “inertia”.

### v2

@@ 3.2 Constrained maximum occupancy
Thus, identifying a lack of action availability with a low-entropy state is plausible only in toy scenarios in which the entropy increase induced within the overall system by the dissolution of the agent is ignored. Death ought to be attractive to MOP agents unless they possess an a priori distinction between agent and environment, i.e. a “sense of self”. The upshot is that the implicit constraint enabling the emergence of goal-directed behavior in such agents is, in the general case, not simply long-term entropy maximization but also the existence of an agent with a repertoire of actions, encoded in the very partitioning of the space into action and state variables. Effectively, this amounts to a version of the “controllability” constraints that appear explicitly in active inference and empowerment, as the agent must exert control sufficient to enable homeostasis (i.e. the maintenance of internal states against dissipative forces). [^13] We may note that the ability to predict the entropy of future states so as to compute state value—however this is implemented—also corresponds to some local disequilibrium and thus imposes a de facto constraint on entropy (here, we are explicitly considering the entropy associated with the internal states of the agent).
@@ 3.3 Model evidence and the will to live
Despite the arguments just given, the inversion of traditional assumptions about the relationship between exploration and reward highlighted by the MOP is appealing, as entropy maximization (albeit under constraints) appears to be an essential feature of intelligence and life (England, 2013; Wissner-Gross and Freer, 2013; Sakthivadivel, 2022; Costa, 2024), more constant across distinct forms of life than any particular reward-seeking behavior. The idea that future path occupancy, as measured by entropy (Ramirez-Ruiz et al., 2022), is tantamount to remaining alive is one way of understanding the place of entropy maximization at the heart of accounts of intrinsic motivation.
We have seen however that in order to reproduce the goal-oriented behavior characteristic of complex biological intelligence, it is necessary to maximize entropy under the constraint that the agent’s existence, operationalized as a conditional independence between internal and external states (Friston, 2019) (which appears in simple models as an action-state partition), is maintained. Taking a page from Schopenhauer (Schopenhauer and Payne, 1958), intrinsic motivation may then be cast as simply the “will to live”, i.e. to persist as a living (moving, changing) thing, a basal impulse that takes different particular forms depending on local constraints (generative models). These constraints shape the primary motivational force of entropy production, such that conditional independence structures are maintained.